# Windowed MAPF with Completeness Guarantees

### v1

@@ 3.2 Reasoning about Coupled Agents
One crucial observation is that it can be non-trivial to determine which agents are coupled once the action generator returns the next state. Agents could be next to each other but still independent, or on the flip side be non-adjacent but coupled. However, instead of reasoning about coupled agents after the action generator, we can use the action generator itself as it must have reasoned about agent interactions to return a valid next action. All modern MAPF heuristic search planners reason about agent interactions internally, e.g. M* explicitly couples agents that intersect (Wagner and Choset 2011), PIBT’s priority inheritance reasons about colliding agents, and CBS resolves conflicts between intersecting agents. Thus, we require that the action generator additionally returns groups of interacting agents. We highlight that this can be generally done with bookkeeping and without much added compute to existing MAPF solvers.
@@
Given a joint configuration $\mathcal{C}$ and a set of HPs for groups of agents at various $\mathcal{C}_{Gr}$, we then compute $h(\mathcal{C})=h_{BD}(\mathcal{C})+\sum_{i}h_{p}(\mathcal{C}_{Gr_{i}})$ for a disjoint set of groups whose locations match the configuration.
@@ 3.3 Overall WinC-MAPF Framework
Thus, we require the AG to incorporate $h(\mathcal{C})$ updates when reasoning about which actions to execute as well as detect the coupled agents for the chosen actions. We focus on optimal windowed action generators that determine the action sequence minimizing $\operatorname*{arg\,min}_{C^{\prime}}c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C})$. We show in our proof in the appendix that these AGs retain completeness guarantees even though they plan partial paths. Future work could show how certain non-optimal windowed AGs could still retain completeness guarantees in our framework.
@@
Given a finite undirected search space and: (1) the initial heuristic is admissible, (2) our AG picks $\mathcal{C}^{\prime}=\operatorname*{arg\,min}c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$ and identifies coupled agents groups, then WinC-MAPF with its grouped update equation is complete (i.e. all agents will eventually reach their goals if a solution exists).
We can now effectively plan in the composite state-space while applying heuristic penalties to groups of agents. Algorithm 1 describes this procedure. We have a set of Heuristic Penalties (line 2) which we pass into the AG with the current configuration (line 4). The AG returns the next configuration (or actions leading to it) as well as a list of agent groups. For each group $Gr$, we compute the update equation with respect to the group’s joint configuration $\mathcal{C}_{Gr}$ and if it is greater than 0, we add the subgroup configuration and penalty into our library of heuristic penalties. Then, we move agents to the picked locations and replan.
@@
4:         $\mathcal{C}^{\prime}$, ListOfGroups = AG($\mathcal{C}^{cur}$, $\mathcal{H}$)
@@
6:              $h_{new}=c(\mathcal{C}^{cur}_{Gr},\mathcal{C}_{Gr}^{\prime})+h(\mathcal{C}_{Gr}^{\prime})$
7:              $penalty=h_{new}-h_{BD}(\mathcal{C}^{cur}_{Gr})$
@@
10:         $\mathcal{C}^{cur}\leftarrow\mathcal{C}^{\prime}$triangleright Move agents
The main assumptions we have currently is that we have a centralized AG and we have a perfect backward Dijkstra heuristic for each agent. The first assumption allows the AG to reason about heuristic penalties between coupled agents. The second assumption is not strictly required but simplifies the problem as we do not need to do single-agent heuristic updates. Both of these assumptions are common in current MAPF literature, and prior work has shown that for certain state-of-the-art methods like LaCAM, a perfect backward Dijkstra heuristic is required (Veerapaneni et al. 2024). Both these assumptions can be relaxed in future work.
The next section describes Single-Step CBS, an action generator that satisfies the two properties required by WinC-MAPF (identifying agent groups and optimally solving a windowed problem).
Figure 2: We depict an illustrative situation where Single-Step CBS needs to determine the best joint configuration to move agents to given heuristic penalties (left). Section 4.1 walks through this figure and describes how naively incorporating heuristic penalties in CBS results in incorrect solutions (middle). Our innovation is to introduce “heuristic conflicts” which allows SS-CBS to find the optimal solution (right).
@@ 4 Single-Step CBS
We want to design a windowed solver that incorporates heuristic penalties and optimally solves $W=1$, i.e. finds the next best step that minimizes $c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$. We note that in MAPF with $N$ agents and an individual action space of size 5, naively computing this requires generating all $5^{N}$ possible neighboring configurations as certain configurations may have heuristic penalties. Thus we employ CBS to intelligently find the optimal single-step configuration.
Since we want a windowed solver with $W=1$, Single-Step CBS (SS-CBS) only considers conflicts within the first timestep. However, regular CBS does not operate in the joint configuration space of MAPF problems and instead exploits the structure of MAPF to iteratively plan agents individually. One key assumption in CBS is that it can minimize the joint $c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$ by minimizing individual $c(s,s^{\prime})+h(s^{\prime})$ subject to constraints. Thus, incorporating HPs which work on the joint configuration of multiple coupled agents breaks this assumption and requires careful reasoning. Our main innovation lies in modifying SS-CBS to return the optimal solution given heuristic penalty updates. A minor additional modification is returning groups of conflicting agents.
@@ 4.1 Handling Heuristic Penalties with Constraints
Incorporating updates to $h$ via heuristic penalties is non-trivial in CBS. A naive way to incorporate a heuristic penalty in CBS is to plan CBS normally and just add the heuristic penalty cost to CT nodes whose configurations match the penalty. However, we show that this fails to find an optimal solution. A second naive way is to incorporate the penalty in the low-level search. We show how this similarly fails. The core conceptual issue with naively incorporating heuristic penalties is that the high-level/low-level is not fully aware of the heuristic penalties until after it has planned paths, so it is unable to avoid them beforehand.
Figure 2 depicts Single-Step CBS in a scenario where two agents want to swap their location from their starting configuration ($R_{1},R_{2}$ want to reach $D,A$ respectively, top left). The left purple box shows two Heuristic Penalties with their corresponding configurations and penalty values (50 and 20 respectively). We created the two HPs for this example; in the real system, HPs would be created from previous iterations of execution and applying the subgroup logic and update equation described in Section 3.2. Given the starting configuration and the two HPs, SS-CBS needs to find the optimal next configuration that minimizes $c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$. Since the cost of all actions is 1, $c(\mathcal{C},\mathcal{C}^{\prime})=2$ regardless of the chosen $\mathcal{C}^{\prime}$ configuration. Thus, we focus on minimizing $h(\mathcal{C}^{\prime})$. In our example, the optimal next configuration is $(R_{1},B),(R_{2},D)$ which has a heuristic of 5.
@@
We take a look at generating the root CT node in our example (middle yellow box). If the root node first plans $R_{2}$, then $R_{2}$ moves to $C$ as this reduces its heuristic (middle box, top row). We do not incur a HP as we do not know the configuration of $R_{1}$ yet. When we plan $R_{1}$, the low-level search has $R_{1}$ minimize its single agent heuristic and move to $B$ which results in the root CT node with $(R_{1},B),(R_{2},C)$ (middle box, middle row). This then incurs HP1’s penalty of 50. Note that there are no agent conflicts and thus CBS will return this solution with a net-heuristic of $4+50=54$. This is substantially worse than the true solution.
@@
Like before, we can plan $(R_{2},C)$ which does not incur any penalty as we do not have $R_{1}$’s position. When planning for $R_{1}$, we know $(R_{2},C)$ and the search will penalize $(R_{1},B)$ by HP1 and instead picks $(R_{1},A)$ which is only penalized by HP2 (middle box, bottom row). This results in a net heuristic of $5+20=25$ which is again not optimal.
Solution: Introducing “Heuristic Conflicts” Our idea is thus not to incorporate the heuristic penalty immediately. Instead, when CBS encounters a joint configuration that would incur a penalty, it marks the CT node with a “heuristic conflict” without adding the penalty into the CT node’s heuristic value yet. Formally, a heuristic conflict occurs when agents’ positions match the positions of a HP. Resolving the heuristic conflict requires applying regular (negative) vertex constraints on each agent in the heuristic conflict which forces them to avoid the joint configuration (and thus penalty) as well as one CT node with positive vertex constraints which requires the agents to be at the penalty joint-configuration and only then incurring the HP111“Negative” vertex constraints avoid vertices while “positive” vertex constraints force agents to certain vertices..
@@
4.2 Detecting Heuristic Agent Groups
SS-CBS should also return groups of agents that are coupled. Our main observation is that agent conflicts directly denote coupled agents. If agents are directly coupled, they must have a conflict between them that got resolved. Similarly, independent agents will not conflict with each other.
There is the possibility of indirect interactions. In particular, $R_{1}$ could conflict with $R_{2}$, causing $R_{2}$ to replan which then conflicts with $R_{3}$. In this case, $R_{1}$’s action of $R_{2}$ directly caused an interaction with $R_{3}$, so $R_{1}$ and $R_{3}$ are indirectly coupled. Thus we can determine disjoint groups of dependent agents by first generating non-disjoint groups of agents for each resolved conflict and then merging groups with shared agents. In our example, we start with $(R_{1},R_{2})$ and $(R_{2},R_{3})$ and will end with $(R_{1},R_{2},R_{3})$ after merging.
We note that we only care about resolved conflicts that occur on CT nodes on the one branch that led to the outputted joint state. Conflicts in other branches do not affect the movements of agents in the outputted joint state. Thus, agents in vertex, edge, or heuristic conflicts on the CT branch to the goal node will be grouped together.
@@ 4.3 Subtleties
A powerful technique for speeding up CBS is Enhanced CBS (Barer et al. 2014) which replaces the low-level and high-level optimal searches with bounded-suboptimal focal searches. Thus, an obvious idea to potentially improve SS-CBS is to introduce these focal searches. As stated in Section 3.3, our proof of completeness only applies to optimal AGs. Hence we set $w_{so}=1$ which results in only using the focal search for tie-breaking. For experimental curiosity, we include results for SS-CBS with a suboptimal high-level search in the appendix (although we do not prove completeness for this suboptimal AG).
We have two additional subtleties. First, there is an important implementation nuance for determining which HP to apply to a configuration if multiple HPs were applicable. Second, we introduced a tiebreaking mechanism that, when faced with two solutions of equal cost, prioritizes reducing the heuristic of some agents over others. Due to space constraints, the details are included in the appendix.
Figure 3: We compare our method SS-CBS (blue), which has $w=1$, against windowed ECBS with different window sizes $w=\{1,2,4,8,16\}$. SS-CBS is theoretically complete (windowed ECBS is not) and also outperforms windowed ECBS empirically.
@@ 5 Experiments
Our experiments demonstrate the empirical performance of our theoretically complete SS-CBS algorithm. We first evaluate SS-CBS on standard benchmark maps (Stern et al. 2019) and observe that SS-CBS is indeed able to outperform the windowed baselines. We then evaluate SS-CBS on high-congestion small maps and showcase SS-CBS’s superiority in this regime. Our appendix contains additional analysis and results. We highlight that there do not exist any complete windowed baselines. We thus compare against windowed ECBS with window $w=\{1,2,4,8,16\}$. For consistency with SS-CBS, we use a suboptimality factor of $w_{so}=1$ for ECBS and no additional CBS improvements.
@@ 5.1 Benchmark Scenarios
Figure 3 shows the results of SS-CBS compared to windowed ECBS without heuristic penalties. We evaluate on 4 benchmark maps (each column: random-32-32-20, ht_chantry, den520d, warehouse-10-20-10-2-1) with a 1 minute timeout across 25 scenarios. We additionally end windowed ECBS in failure when it repeats previously visited configurations 100 times (i.e. deadlock or livelock).
The first row shows how SS-CBS (blue) has an almost strictly better performance than the windowed ECBS baselines. We first highlight that the performance of Windowed ECBS depends significantly on the map, with no single window value dominating. Windowed ECBS with $w=1,2$ sometimes perform poorly due to their small window size which leads to deadlock/livelock, while $w=8,16$ sometimes suffer from large runtimes. SS-CBS is able to consistently perform better than the corresponding windowed ECBS method even though SS-CBS plans only a single step. SS-CBS’s performance on warehouse-10-20-2-1 shows how it performs well in scenarios that require long-term planning (as shown by Windowed ECBS requiring a larger window size to have non-trivial success rate).
The second row shows the per-iteration runtime of each method, with the median runtime in solid and the maximum (longest) iteration runtime in dashed. The difference between the median and maximum highlights how congested iterations can take a significant amount of time (e.g. 10s of seconds for one optimal step) compared to typical iterations. We observe how SS-CBS’s runtime is usually in line with ECBS $w=1,2,4$ and substantially smaller than $w=8,16$.
The solution cost of SS-CBS was roughly 2-8% higher than windowed ECBS. We observed that across all maps, the number of HPs created by SS-CBS grew roughly linearly along with the number of agents, but the number of HPs encountered in the CT grew exponentially.

### v2

@@ 3.2 Disjoint Agent Groups
Our objective is, given $\mathcal{C},\mathcal{C}^{W}$, to detect these groups of agents and apply heuristic penalties to just the group of agents rather than the entire configuration. One crucial observation is that it can be non-trivial to determine which agents are coupled once the AG returns the next state. Agents could be next to each other but not block each other, or on the flip side be non-adjacent but coupled. However, instead of reasoning about coupled agents after the AG, we can leverage the AG itself as it must have internally reasoned about agent interactions to return a valid next action. All modern MAPF heuristic search planners reason about agent interactions internally, e.g., M* explicitly couples agents that intersect (Wagner and Choset 2011), PIBT’s priority inheritance reasons about colliding agents, and CBS resolves collisions between intersecting agents. Thus, we require that the AG additionally returns groups of interacting agents. We highlight that this can be generally done with bookkeeping and without much added compute to existing MAPF solvers.
@@
Given a configuration $\mathcal{C}$ and a set of HPs for groups of agents at various $\mathcal{C}_{Gr}$, we then compute $h(\mathcal{C})=h_{BD}(\mathcal{C})+\sum_{i}h_{p}(\mathcal{C}_{Gr_{i}})$ for a disjoint set of groups $\{\mathcal{C}_{Gr_{i}}\}$ whose locations match the configuration.
We lastly note that Disjoint Agents Groups is related to Independence Detection in Operator Decomposition (Standley 2010) which dynamically constructs independent subproblems by checking if agents block each other.
@@ 3.3 Overall WinC-MAPF Framework
Altogether, our framework requires an Action Generator AG with the following two properties: (1) finds $\operatorname*{arg\,min}_{\mathcal{C}^{W}}c(\mathcal{C},\mathcal{C}^{W})+h(\mathcal{C}^{W})$, and (2) computes disjoint agent groups for $\mathcal{C}\rightarrow\mathcal{C}^{W}$.
Given such an action generator, we are guaranteed to eventually reach the goal if a solution exists via Algorithm 1. We have a set of Heuristic Penalties (line 2) which we pass into the AG with the current configuration (line 4). The AG returns a partial path corresponding to configuration $\mathcal{C}^{W}$ as well as a list of disjoint agent groups. For each group $Gr$, we compute the update equation with respect to the group’s configuration $\mathcal{C}_{Gr}$ and if it is greater than 0, we add the group configuration and penalty into our library of heuristic penalties. Then, we move agents and repeat.
@@
Given a finite bidirectional graph and: (1) an initial Backward Dijkstra heuristic, (2) our AG picks $\operatorname*{arg\,min}_{\mathcal{C}^{W}}c(\mathcal{C},\mathcal{C}^{W})+h(\mathcal{C}^{W})$ and identifies disjoint agents groups, then WinC-MAPF with its update equation (Eq. 1) applied on group configurations is complete, i.e. all agents will eventually reach their goals if a solution exists.
Section B.2 contains a formal proof. Briefly, the key is showing how $h(\mathcal{C})$ remains admissible even with the grouped HPs, which we do by reasoning about the interaction of agents inside of and between disjoint agent groups.
@@
4:         $\mathcal{C}^{W}$, ListOfGroups = AG($\mathcal{C}^{cur}$, $\mathcal{H}$)
@@
6:              $h_{n}\leftarrow U(\mathcal{C}^{cur}_{Gr},\mathcal{C}_{Gr}^{W})$triangleright Equation (1)
7:              $penalty\leftarrow h_{n}-h_{BD}(\mathcal{C}_{Gr}^{W})$
@@
10:         $\mathcal{C}^{cur}\leftarrow C^{0<i\leq W}$triangleright Move at least one step
The main assumptions we have currently are that we have a centralized AG and we have a perfect backward Dijkstra heuristic for each agent. The first assumption allows the AG to reason about heuristic penalties between coupled agents. The second assumption is not strictly required but simplifies the problem as we do not need to do single-agent heuristic updates. Both of these assumptions are common in current MAPF literature, and prior work has shown that for certain state-of-the-art methods like LaCAM, a perfect backward Dijkstra heuristic is required (Veerapaneni et al. 2024). Both of these assumptions can be relaxed in future work.
Section 4 describes Single-Step CBS, an action generator that satisfies the two properties required by WinC-MAPF.
Figure 2: We depict an illustrative situation where SS-CBS needs to determine the best next configuration given heuristic penalties (HP, left). Section 4.1 describes this figure and how naively incorporating HPs in CBS results in incorrect solutions (middle). Our innovation is to introduce “heuristic conflicts” which allows SS-CBS to find the optimal solution (right).
@@ 4 Single-Step CBS
We want to design a windowed solver that incorporates heuristic penalties and optimally solves $W=1$, i.e. finds the next best step that minimizes $c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$. We focus on solving $W=1$ as this is still a challenging problem. In MAPF with $N$ agents and an individual action space of size 5, naively computing the optimal action requires generating all $5^{N}$ possible neighboring configurations as heuristic penalties could arbitrarily be placed! Thus we employ CBS [^1] Readers unfamiliar with CBS should read Sharon et al. (2015), an optimal full horizon planner, to intelligently find the optimal single-step configuration.
Since we want a windowed solver with $W=1$, Single-Step CBS (SS-CBS) only considers conflicts within the first timestep. However, regular CBS does not operate in the configuration space of MAPF problems and instead exploits the structure of MAPF to iteratively plan agents individually. One key assumption in CBS is that it can minimize the joint $c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$ by minimizing individual $c(s,s^{\prime})+h(s^{\prime})$ subject to constraints. Thus, incorporating HPs which work on the configuration of disjoint agent groups breaks this assumption and requires careful reasoning. Our main innovation lies in modifying SS-CBS to return the optimal solution given heuristic penalty updates. A minor additional modification is returning disjoint agent groups.
@@ 4.1 Handling Heuristic Penalties with Constraints
Incorporating updates to $h$ via heuristic penalties is non-trivial in CBS. A naive way to incorporate a HP in CBS is to plan CBS normally and just add the HP cost to Constraint Tree (CT) nodes whose configurations match the penalty. However, we show that this fails to find an optimal solution. A second naive way is to incorporate the penalty in the low-level search which similarly fails. The core conceptual issue with naively incorporating HPs is that the high-level/low-level is not fully aware of the HPs until after it has planned paths, so it is unable to avoid them beforehand.
Figure 2 depicts Single-Step CBS in a scenario where two agents want to swap their locations from their starting configurations (i.e., $R_{1},R_{2}$ want to reach $D,A$ respectively, top left). The left purple box shows two HPs with their corresponding configurations and penalty values (50 and 20 respectively). We created the two HPs for this example; in the real system, HPs would be created from previous iterations of execution and applying the subgroup logic and Equation 1 described in Section 3.2. Given the starting configuration and the two HPs, SS-CBS needs to find the optimal next configuration that minimizes $c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$. Since the cost of all actions is 1, $c(\mathcal{C},\mathcal{C}^{\prime})=2$ regardless of the chosen $\mathcal{C}^{\prime}$ configuration. Thus, we focus on minimizing $h(\mathcal{C}^{\prime})$. In our example, the optimal next configuration is $(R_{1},B),(R_{2},D)$ which has a heuristic of 5.
@@
We take a look at generating the root CT node in our example (middle yellow box). If the root node first plans $R_{2}$, then $R_{2}$ moves to $C$ as this reduces its heuristic (middle box, top row). We do not incur a HP as we do not know the configuration of $R_{1}$ yet. When we plan $R_{1}$, the low-level search has $R_{1}$ minimize its single agent heuristic and move to $B$ which results in the root CT node with $(R_{1},B),(R_{2},C)$ (middle box, middle row). This then incurs HP1’s penalty of 50. Note that there are no agent collisions and thus CBS will return this solution with a net-heuristic of $4+50=54$. This is substantially worse than the true solution.
@@
Like before, we can plan $(R_{2},C)$ which does not incur any penalty as we do not have $R_{1}$’s location. When planning for $R_{1}$, we know $(R_{2},C)$ and the search will penalize $(R_{1},B)$ by HP1 and instead picks $(R_{1},A)$ which is only penalized by HP2 (middle box, bottom row). This results in a net heuristic of $5+20=25$ which is again not optimal.
Solution: Introducing “Heuristic Conflicts” Our idea is thus not to incorporate the heuristic penalty immediately. Instead, when CBS encounters a configuration that would incur a penalty, it marks the CT node with a “heuristic conflict” without adding the penalty into the CT node’s heuristic value yet. Formally, a heuristic conflict occurs when agents’ locations match the configuration of a HP. Resolving the heuristic conflict requires applying regular (negative) vertex constraints on each agent in the heuristic conflict which forces them to avoid the configuration (and thus penalty) as well as one CT node with positive vertex constraints which requires the agents to be at the penalty configuration and only then incurring the HP222“Negative” vertex constraints avoid vertices while “positive” vertex constraints force agents to certain vertices..
@@
4.2 Detecting Disjoint Agent Groups
SS-CBS should also return disjoint agent groups where each disjoint group contains coupled agents. Our key observation is that coupled agents must have conflict(s) between them that got resolved. Similarly, independent agents will not conflict with each other.
One small subtly; coupled agents must have conflicts, but agents with conflicts may not be coupled, e.g. an agent could tie-break poorly and have an “unnecessary” conflict. From our definition of disjoint agent groups, having extra agents in groups is acceptable.
There is the possibility of indirect interactions. In particular, $R_{1}$ could conflict with $R_{2}$, causing $R_{2}$ to replan which then conflicts with $R_{3}$. In this case, $R_{1}$’s action of $R_{2}$ directly caused an interaction with $R_{3}$, so $R_{1}$ and $R_{3}$ are indirectly coupled. Thus we can determine disjoint groups of coupled agents by first generating non-disjoint groups of agents for each resolved conflict and then merging groups with shared agents. In our example, we start with $(R_{1},R_{2})$ and $(R_{2},R_{3})$ and will end with $(R_{1},R_{2},R_{3})$ after merging.
We note that we only care about resolved conflicts that occur on CT nodes on the solution branch that led to the outputted configuration. Thus when planning, once a goal CT node is reached, we backtrack and store all resolved vertex, edge, and heuristic conflicts as individual groups, and then merge groups together to get our disjoint agent groups.
@@ 4.3 Subtleties
We have two additional subtleties. First, there is an important implementation nuance for determining which HP to apply to a configuration if multiple HPs were applicable. Second, we introduced a tiebreaking mechanism that, when faced with two solutions of equal cost, prioritizes reducing the heuristic of some agents over others. Both are discussed in detail in the appendix.
Figure 3: We compare our method SS-CBS (blue), which has $W=1$, against windowed CBS (wCBS) with different window sizes $W=\{1,2,4,8,16\}$. SS-CBS is theoretically complete (wCBS is not) and also outperforms wCBS empirically.
@@ 5 Experiments
Our experiments demonstrate the empirical performance of our theoretically complete SS-CBS algorithm. We first evaluate SS-CBS on standard benchmark maps (Stern et al. 2019) and observe that SS-CBS is indeed able to outperform the windowed baselines. We then evaluate SS-CBS on high-congestion small maps and showcase SS-CBS’s superiority in this regime. Our appendix contains additional analysis and results. We highlight that there do not exist any complete windowed baselines. We thus compare against CBS with window $W=\{1,2,4,8,16\}$ (wCBS). All algorithms were implemented in C++ and run on a PC with a 2.30 GHz Intel i7-11800K CPU.
@@ 5.1 Benchmark Scenarios
Figure 3 shows the results of SS-CBS compared to wCBS without heuristic penalties. We evaluate on 4 benchmark maps (each column: random-32-32-20, ht_chantry, den520d, warehouse-10-20-10-2-1) with a 1 minute timeout across 25 scenarios. We additionally end wCBS in failure when it repeats previously visited configurations 100 times (i.e. deadlock or livelock).
The first row shows how SS-CBS (blue) has an almost strictly better performance than the wCBS baselines. We first highlight that the performance of wCBS depends significantly on the map, with no single window value dominating. wCBS with $W=1,2$ sometimes perform poorly due to their small window size which leads to deadlock/livelock, while $W=8,16$ sometimes suffer from large runtimes. SS-CBS is able to consistently perform better than the corresponding wCBS method even though SS-CBS plans only a single step. SS-CBS’s performance on warehouse-10-20-2-1 shows how it performs well in scenarios that require long-term planning (as shown by wCBS requiring a larger window size to have non-trivial success rate).
The second row shows the per-iteration runtime of each method, with the median runtime in solid and the maximum (longest) iteration runtime in dashed. Limitation: The difference between the median and maximum highlights how congested iterations can take a significant amount of time (e.g. 10s of seconds for one optimal step) compared to typical iterations. All the failure instances for SS-CBS occurred due to the large runtime of those bottleneck iterations. We observe how SS-CBS’s runtime is usually in line with wCBS $W=1,2,4$ and substantially smaller than $W=8,16$.
The solution cost of SS-CBS is roughly 2-8% higher than wCBS due to SS-CBS’s myopic planning window. We observed that the number of HPs created by SS-CBS grew roughly linearly along with the number of agents, but the number of HPs encountered in the CT grew exponentially.